OpenStudy (anonymous):

Answer this for me, Please and Thank you! >.< A 36-foot long fence is placed around a square garden. What is the length of one side of the garden? 4 feet 6 feet 9 feet 12 feet

OpenStudy (anonymous):

perimeter =s^2 36=s^2 square root of 36=6

OpenStudy (anonymous):

A square has 4 equal sides. So you would divide 36 by 4 to get??

OpenStudy (anonymous):

@JaneDoe it would be 9 because 9 x 4 is 36

OpenStudy (anonymous):

36/4 = 9

OpenStudy (anonymous):

omg.. sorry .. i was doing area

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Hahaa its all good

OpenStudy (anonymous):

ty :D

OpenStudy (anonymous):

@EMINEM i need help with a question

OpenStudy (anonymous):

mk

OpenStudy (anonymous):

Its posted on the board

OpenStudy (anonymous):

what do you need?..